# Flavor Equilibration of Supernova Neutrinos: Exploring the Dynamics of Slow Modes

### v1

@@ I Introduction
Over almost the past decade, research on neutrino FCs has focused mainly on understanding fast flavor conversions (FFCs) first discovered in [33, 34, 35] (see [12] for a recent review and references therein). However, the study of neutrino flavor dynamics in neutrino-dense media originally began with what are now called slow flavor conversions (SFCs) [7, 8, 36, 37, 9]. An important distinction between FFCs and SFCs is that the former can occur in the limit of vanishing vacuum oscillation frequency. Unlike FFCs, which occur only when there is a crossing in the angular distribution of electron neutrinos and antineutrinos, e.g. the electron neutrino lepton number (ELN) angular distribution (assuming heavy-lepton neutrinos and antineutrinos have similar distributions) [38, 39, 40], SFCs can exist throughout the SN environment for regions outside the proto-neutron star after the neutronization burst phase when spectra crossings are expected. Despite significant progress in understanding FFCs and their dynamical evolution in realistic CCSN environments [41, 42, 28, 43, 44, 45], SFCs remain considerably underdeveloped. This is primarily because just as the impact of broken spatial and temporal symmetries on the dynamics of SFCs in SNe began to be explored [46, 47, 48, 49], FFCs were discovered. Due to the very short time and length scales of FFCs and their potential presence near neutrino decoupling regions, research in the field quickly shifted toward studying FFCs, leaving SFCs largely overlooked in the literature ever since.
@@
SFCs have recently regained attention, in part due to new insights into the impact of FCs in low-density regions ($\rho\sim 10^{9}-10^{10}\ {\rm g/cm^{3}}$) on SN dynamics [29, 30, 32]. For instance, Refs. [50, 51] recently revisited the linear stability analysis of SFCs and explored different regimes of FCs. While FFCs may occur at various regions across a wide range of densities in the SN environment [52, 53], e.g., a thin shell inside the proto-neutron star [54], the decoupling regions where local $\bar{\nu}_{e}$ emission dominates due to the multidimensional effect [55, 56, 57, 58], and the low-density, neutrino free-streaming region due to scatterings [59, 57], SFCs may become more relevant in low-density regions. This is because at such lower densities, FFCs are mostly not expected to be significant and have been suggested to be unlikely to result in significant flavor conversions [60]. Besides FFCs, the collisional flavor instability (CFI) [61, 62, 63, 64, 65] is expected to exist mainly at decoupling densities of $\rho\sim 10^{10}-10^{12}\,{\rm g/cm^{3}}$ [61, 66, 67]. At low densities, even if CFIs exist, they are expected to be subdominant compared to SFCs. This is because CFIs occur on collisional scales, which are typically $\gtrsim$ a few tens of kilometers in lower-density regions—significantly larger than the characteristic scales of SFCs, which are of the order of $\sim\omega^{-1}=0.7\,\mathrm{km}\times(E/10\,\mathrm{MeV})$. Here, $\omega=\Delta m^{2}/2E$, where $\Delta m^{2}$ denotes the atmospheric neutrino mass-squared difference and $E$ represents the neutrino energy. Nevertheless, the interplay between SFCs and CFIs or FFCs may give rise to nontrivial FC patterns.
@@
In this study, we investigate neutrino flavor equilibration in the context of SFCs for a monochromatic neutrino gas, considering moderate lepton number asymmetries, e.g., $0.7\lesssim n_{\bar{\nu}_{e}}/n_{\nu_{e}}\lesssim 1.3$ without considering momentum changing collisions. To this end, we analyze the evolution of a neutrino gas in periodic boxes and examine the spatially averaged quantities within the box. This approach closely follows the methodology used in the study of flavor equilibration in FFCs [68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81]. Our motivation for taking this approach for SFCs stems from the fact that, in the regions of interest where $\rho\sim 10^{9}-10^{10}\ {\rm g/cm^{3}}$, the collisional scales are expected to be significantly larger than the characteristic scales of SFCs, as discussed above. Under such conditions, slow modes can effectively be treated as fast compared to the collisional scale, justifying our approach. This consideration is particularly important, as the primary goal of this study is to develop a better understanding of the local quasi-equilibrium state of SFCs. This understanding could help develop practical methods to include the effects of FCs in CCSN simulations. [82, 83, 43].
@@ II Quantum Kinetic Equations
$\rho=\begin{pmatrix}\rho_{ee}&\rho_{ex}\\ \rho_{ex}^{*}&\rho_{xx}\end{pmatrix},\hskip 8.53581pt\bar{\rho}=\begin{pmatrix}\bar{\rho}_{ee}&\bar{\rho}_{ex}\\ \bar{\rho}_{ex}^{*}&\bar{\rho}_{xx}\end{pmatrix}\ ,$ (1)
@@
$H_{\rm vac}=\frac{\omega}{2}\begin{pmatrix}-\cos{2\theta_{V}}&\sin{2\theta_{V}}\\ \sin{2\theta_{V}}&\cos{2\theta_{V}}\end{pmatrix}\ ,$ (3)
@@
$H_{\rm mat}=\begin{pmatrix}\lambda&0\\ 0&0\end{pmatrix}\ ,$ (4)
@@
For practical purposes, here we ignore the constant matter contribution to the Hamiltonian, given the fact that it can be rotated away by frame transformation. We then use the matter-suppressed mixing angle [86, 45], which is fixed to be $10^{-5}$ radians. We have checked that our results remain qualitatively unchanged once the matter Hamiltonian is properly included with $\lambda=2\mu$ and vacuum mixing angle of $\theta_{V}=0.11$ radians. Moreover, since $\mu$ and $\omega$ are the only energy scales in the system, we express time and distance in units of $\mu^{-1}$. Consequently, specifying the ratio $\omega/\mu$ allows us to fully determine the flavor evolution of a particular system. For the remainder of the paper, we fix $\mu=1$ so that all quantities are shown as dimensionless, with the understanding that their physical units are scaled by the correct power of $\mu$.
@@ III Simulation of SFC
Figure 1: Initial conditions for the matrix elements of $\rho(z,v_{z})$ and $\bar{\rho}(z,v_{z})$ using $\alpha=0.925$ and $\sigma=0.6$. Upper panel: Spatial dependence of the real part of the off-diagonal term, ${\rm Re}(\rho_{ex})$, for various angular bins in range $-1\leq v_{z}\leq 1$. The spatial distribution is localized around $z=0$ for $-50\leq z\leq 50$. Lower panel: Angular distributions of the diagonal terms of the matrix elements at $z=0$. There is no angular crossings between the electron and anti-electron distributions. The non-electron flavors $\rho_{xx}$ and $\bar{\rho}_{xx}$ are $\sim 10^{-5}$ at the most forward bin with $v_{z}=1$.
@@
where $\sigma$ is the width of a half-Gaussian distribution and $A^{-1}=\sigma\sqrt{\pi/2}\ {\rm erf}(\sqrt{2}/\sigma)$ with ${\rm erf}$ the error function. For antineutrinos, we assume their angular distribution follows $\bar{g}(v_{z},\sigma)=\alpha g(v_{z},\sigma)$ with $\alpha=n_{\bar{\nu}_{e}}/n_{\nu_{e}}$ being the neutrino-antineutrino asymmetry factor. The choice of the same width parameter $\sigma$ for neutrinos and antineutrinos is to avoid appearance of ELN crossings that can trigger FFCs.
@@ IV Linear Stability Analysis
$\rho$ $=$ $\begin{pmatrix}\rho_{ee}(t_{0})-\Delta(t)&\rho_{ex}(t)\\ \rho_{ex}^{*}(t)&\rho_{xx}(t_{0})+\Delta(t)\end{pmatrix}\ ,$ (13)
$\bar{\rho}$ $=$ $\begin{pmatrix}\bar{\rho}_{ee}(t_{0})-\bar{\Delta}(t)&\bar{\rho}_{ex}(t)\\ \bar{\rho}_{ex}^{*}(t)&\bar{\rho}_{xx}(t_{0})+\bar{\Delta}(t)\end{pmatrix}\ ,$ (14)
@@
isotropic, NO
isotropic, IO
anisotropic, NO
anisotropic, IO
Figure 4: Distributions (upper), absolute values of the eigenvectors $\vec{V}$ (middle), and a simulation snapshot of $\langle P_{3}\rangle$ and $\langle\bar{P}_{3}\rangle$ in linear regime (bottom). The columns correspond to isotropic angular distributions ($\sigma\rightarrow\infty$, $k=0$) for NO and IO, and anisotropic angular distributions ($\sigma=0.6$, $k=0.075$) for NO and IO, respectively. We keep the parameters $\alpha=0.85$, and $|\omega|=0.01$ fixed. The normalization is such that $\sum|{\vec{V}}|^{2}=1$. In the upper panels, the color distinguishes the angular bins from $-1$ (red) to $1$ (purple), while thick and thin markers correspond to $\nu_{e}$ and $\bar{\nu}_{e}$, respectively. The bottom panels show the angular distributions of the averaged $P_{3}$ and $\bar{P}_{3}$ in the linear regime ($t=60$ for NO, and $t=20$ for IO). The angular modes at which $P_{3},\bar{P}_{3}$ are closest to 1 correspond to the locations where $|{\vec{V}}|$ has a minimum. The shape of the angular distribution in the linear regime is described by the shape of $|{\vec{V}}|$.
@@
where $\delta_{ij}$ is the Kronecker delta which accounts for the extra contribution to the diagonal terms. The matrix $\Lambda$ is completely determined by the initial conditions of the system and does not depend on the dynamical variable $\rho_{ex}$. We compute the eigenvalues of $\Lambda$ and search for the eigenvalue with the largest positive imaginary component. The imaginary part of this $\Omega$ represents the growth rate of the unstable solution, or in other words, $\vec{V}\sim e^{{\rm Im}{(\Omega)}t}$ grows at a rate fixed by Im$(\Omega)$.
In Fig. 2, we show the dispersion relation of a system with $\omega=\pm 0.01$ and (anti)neutrino angular distributions defined by Eqs. 8 and 10 with $\alpha=0.850$ and $\sigma=0.6$. The NO case (blue) generically shows smaller growth rates than the IO case (red) across the $k$-domain. Moreover, the IO case has a wider range of instability in $k$ than the NO case. Both mass orderings have a global maximum around the $k=0.075$ mode which is marked by the dotted line. Additionally, one can see that the IO case is approximately dominated by the $k=0$ mode, while the NO case has a subdominant homogeneous mode and is thus driven by $k\neq 0$ modes. To paint a more general picture of the instability region, in Fig. 3 we show the growth rates of the parameter space comprised by $0.8\leq\alpha\leq 1.2$ and $-0.02\leq\omega\leq 0.02$. In the top panel of Fig. 3, one can see that for the NO case, the systems are more strongly dominated by inhomogeneous ($k\neq 0$) modes when $\alpha$ deviates from 1 as previously discussed. Moreover, the growth rates are approximately symmetric around $\alpha=1$, e.g., the growth rates of mode $k$ for $\alpha=0.9$ is roughly the same as that of mode $-k$ for $\alpha=1.1$. The bottom panel of Fig. 3 shows the instability region for the dominant $k$ mode for any given pair of parameters $(\omega,\alpha)$. One can see that the $\omega<0$ (IO) quadrant has generally larger growth rates that span a wider range of values of $\alpha$ than $\omega>0$ (NO).
@@ V Flavor Equilibration of SN Neutrinos Induced by Slow Modes
As illustrated in Fig. 5, the averaged survival probabilities for neutrinos and antineutrinos reach equilibration after a few $\omega^{-1}$ in both neutrino mass orderings. While the antineutrinos approach a state closer to equipartition, the neutrinos settle at a value consistent with the conservation of the total (E–X)LN (which is guaranteed to a good degree given the suppressed effective vacuum mixing angle; see Appendix A). Due to the approximate conservation of (E–X)LN, the final flavor conversion of the neutrinos is related to that of the antineutrinos through
@@
Figure 9: Time evolution of the $P_{3}(t,z_{,}v_{z})$ component of the neutrino polarization vector in the domain box spanned by the spatial variable $-50\leq z\leq 50$ (horizontal axis) and the angular variable $-1\leq v_{z}\leq 1$ (vertical axis). The evolution corresponds to the benchmark system with $\alpha=0.925$ and $\sigma=0.6$. The panels in the left column correspond to snapshots of the flavor evolution for the $\omega=0.01$ (NO) case, while the right panels to the $\omega=-0.01$ (IO) case.
In almost all cases, antineutrinos in the IO tend to evolve toward a state closer to perfect equipartition. While there is no obvious explanation for this difference, we speculate that this behavior is associated with the distinct instability characteristics of the neutrino gas in the linear regime for NO and IO. In particular, as illustrated in Fig. 2, in IO, a wider range of Fourier modes is unstable in the linear regime (and with higher growth rate), implying that the initial evolution of inhomogeneity-induced decoherence could proceed more efficiently in IO, with the homogeneous mode ($k=0$) being almost the dominantly unstable mode. This fundamental difference between NO and IO is also associated with the different shapes of the eigenvectors of the unstable modes, as illustrated in Fig. 4. While in the IO case, all angle bins are almost equally affected by instabilities, in the NO case, the impact varies depending on the angle.
@@
While near flavor equipartition was also observed to occur due to the multi-angle-induced decoherence previously reported in 1D models for $\alpha$ values very close to 1 for IO [91], our key finding is that in the presence of inhomogeneities, this becomes a more generic outcome for the neutrino gas. As shown in Fig. 10, the neutrino gas reaches near flavor equipartition for $\alpha=1$ cases without advection in IO (consistent with what was observed in homogeneous models). Near equipartition for all other cases occurs only when advection is included in the model. This implies that the equilibration unearthed in this study can be considered as an inhomogeneity-induced decoherence rather than a multi-angle-induced decoherence discovered previously in the literature.
@@
So far, we have only discussed cases where $\alpha\leq 1$. A very similar flavor equilibration outcome can occur for $\alpha>1$, as shown in Fig. 11. However, there is an important yet predictable difference in this scenario: the roles of neutrinos and antineutrinos are reversed. Specifically, for $\alpha>1$, it is the neutrinos that reach near flavor equipartition, while the antineutrinos adjust to a value that ensures the approximate (E-X)LN conservation. In addition and similar to $\alpha<1$, neutrinos reach a state closer to equipartition for IO.

### v2

@@ I Introduction
Over almost the past decade, research on neutrino FCs has focused mainly on understanding fast flavor conversions (FFCs) first discovered in [33, 34, 35] (see [12] for a recent review and references therein). However, the study of neutrino flavor dynamics in neutrino-dense media originally began with what are now called slow flavor conversions (SFCs) [7, 8, 36, 37, 9]. An important distinction between FFCs and SFCs is that the former can occur in the limit of vanishing vacuum oscillation frequency. Unlike FFCs, which occur only when there is a crossing in the angular distribution of electron neutrinos and antineutrinos, e.g. the electron neutrino lepton number (ELN) angular distribution (assuming heavy-lepton neutrinos and antineutrinos have similar distributions) [38, 39, 40], SFCs can exist throughout the SN environment for regions outside the proto-neutron star (PNS) after the neutronization burst phase when spectra crossings are expected. Despite significant progress in understanding FFCs and their dynamical evolution in realistic CCSN environments [41, 42, 28, 43, 44, 45], SFCs remain considerably underdeveloped. This is mainly because, just as the impact of broken spatial and temporal symmetries on the dynamics of SFCs in SNe began to be explored [46, 47, 48, 49], FFCs were discovered. Due to the very short time and length scales of FFCs and their potential presence near neutrino decoupling regions, research in the field quickly shifted toward studying FFCs, leaving SFCs largely overlooked in the literature ever since.
@@
SFCs have recently regained attention, in part due to new insights into the potential impact of FCs in low-density regions ($\rho\sim 10^{9}-10^{10}\ {\rm g/cm^{3}}$) on SN dynamics [29, 30, 32]. For instance, the linear stability analysis based on a spherically symmetric SN model performed in Ref. [44] found instabilities associated with SFCs exists at similar densities behind the shock during the accretion phase. Refs. [50, 51] recently also revisited the linear stability analysis of SFCs and explored different regimes of FCs. These analyses quantify that SFC typically exhibit inverse growth rates of the order of order of $\sim\omega^{-1}=0.7\,\mathrm{km}\times(E/10\,\mathrm{MeV})$, where $\omega=\Delta m^{2}/2E$, $\Delta m^{2}$ denotes the atmospheric neutrino mass-squared difference, and $E$ represents the neutrino energy. Although the inverse growth scale of SFC is of astrophysical size, the unstable region is typically broad enough so that they may not be suppressed in the SN environment. While FFCs may occur at various regions across a wide range of densities in the SN environment [52, 53], e.g., a thin shell inside the PNS [54], the decoupling regions where local $\bar{\nu}_{e}$ emission dominates due to the multidimensional effect [55, 56, 57, 58], and the low-density, neutrino free-streaming region due to scatterings [59, 57], SFCs may become more relevant in low-density regions. This is because at such lower densities, FFCs are mostly not expected to be significant and have been suggested to be unlikely to result in significant flavor conversions [60]. Besides FFCs, the collisional flavor instability (CFI) [61, 62, 63, 64, 65] is expected to exist mainly at decoupling densities of $\rho\sim 10^{10}-10^{12}\,{\rm g/cm^{3}}$ [61, 66, 67]. At low densities, even if CFIs exist, they are expected to be subdominant compared to SFCs. This is because CFIs occur on collisional scales, which are typically $\gtrsim$ a few tens of kilometers in lower-density regions—significantly larger than the SFC characteristic scales of order $\omega^{-1}$. All these indicate that SFC can be the dominant FC of neutrinos in the low-density regions behind the SN shock and can potentially affect SN dynamics, which calls for better understanding of SFC. Moreover, SFC are also expected to occur during the PNS cooling phase and can significantly influence SN nucleosynthesis and neutrino signals.
@@
Early pioneering studies of SFC relied on the assumption of a stationary neutrino gas in the SN environment [7, 8, 36, 37, 9]. However, to fully capture the potential impact of SFC in low-density regions, it is essential to account for both spatial and temporal evolution of the neutrino flavor field. While a few works have performed numerical simulations to study the evolution of different FCs, including SFCs, over static global SN profiles [68, 44], they have adopted artificially quenched neutrino-neutrino interaction potentials, which can affect the outcome of true SFC evolution. In this study, we take a different approach aiming to investigate neutrino flavor equilibration in the context of SFCs for a monochromatic neutrino gas, considering moderate lepton number asymmetries, e.g., $0.7\lesssim n_{\bar{\nu}_{e}}/n_{\nu_{e}}\lesssim 1.3$ without considering momentum changing collisions. To this end, we analyze the evolution of a neutrino gas in periodic boxes and examine the spatially averaged quantities within the box. This approach closely follows the methodology used in the study of flavor equilibration in FFCs [69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82]. Our motivation for taking this approach for SFCs stems from the fact that, in the regions of interest where $\rho\sim 10^{9}-10^{10}\ {\rm g/cm^{3}}$, the collisional scales are expected to be significantly larger than the characteristic scales of SFCs, as discussed above. Under such conditions, slow modes can effectively be treated as fast compared to the collisional scale, justifying our approach. This consideration is particularly important, as the primary goal of this study is to develop a better understanding of the local quasi-equilibrium state of SFCs. We note that although unlike FFCs, the local SFC quasi-equilibrium state cannot be directly used for instantaneous flavor redistribution in global transport or in CCSN simulations, its identification could help develop practical methods to include the effects of FCs in CCSN simulations [83, 84, 43].
@@ II Quantum Kinetic Equations
$\rho=\begin{pmatrix}\rho_{ee}&\rho_{ex}\\ \rho_{ex}^{*}&\rho_{xx}\end{pmatrix},\hskip 8.53581pt\bar{\rho}=\begin{pmatrix}\bar{\rho}_{ee}&\bar{\rho}_{ex}\\ \bar{\rho}_{ex}^{*}&\bar{\rho}_{xx}\end{pmatrix}\$, (1)
@@
$H_{\rm vac}=\frac{\omega}{2}\begin{pmatrix}-\cos{2\theta_{V}}&\sin{2\theta_{V}}\\ \sin{2\theta_{V}}&\cos{2\theta_{V}}\end{pmatrix}\$, (3)
@@
$H_{\rm mat}=\begin{pmatrix}\lambda&0\\ 0&0\end{pmatrix}\$, (4)
@@
For practical purposes, here we ignore the constant matter contribution to the Hamiltonian, given the fact that it can be rotated away by frame transformation. We then use the matter-suppressed mixing angle [87, 45], which is fixed to be $10^{-5}$ radians. We have checked that our results remain qualitatively unchanged once the matter Hamiltonian is included with a representative value of $\lambda=2\mu$ and vacuum mixing angle of $\theta_{V}=0.11$ radians. In realistic CCSN environments, $\lambda/\mu$ can be as large as $\sim 10$ behind the shock during the accretion phase [88, 86] or as small as $\sim 0.1$ between $\sim 50$ and 500 km in the neutrino-driven winds [22, 86].
Moreover, since $\mu$ and $\omega$ are the only energy scales in the system, we express time and distance in units of $\mu^{-1}$. Consequently, specifying the ratio $\omega/\mu$ allows us to fully determine the flavor evolution of a particular system. For the remainder of the paper, we fix $\mu=1$ so that all quantities are shown as dimensionless, with the understanding that their physical units are scaled by the correct power of $\mu$.
@@ III Simulation of SFC
Figure 1: Initial conditions for the matrix elements of $\rho(z,v_{z})$ and $\bar{\rho}(z,v_{z})$ using $\alpha=0.925$ and $\sigma=0.6$. Upper panel: Spatial dependence of the real part of the off-diagonal term, ${\rm Re}(\rho_{ex})$, for various angular bins in range $1\leq v_{z}\leq 1$. The spatial distribution is localized around $z=0$ for $50\leq z\leq 50$. Lower panel: Angular distributions of the diagonal terms of the matrix elements at $z=0$. There is no angular crossings between the electron and anti-electron distributions. The non-electron flavors $\rho_{xx}$ and $\bar{\rho}_{xx}$ are $\sim 10^{-5}$ at the most forward bin with $v_{z}=1$.
@@
where $\sigma$ is the width of a half-Gaussian distribution and $A^{-1}=\sigma\sqrt{\pi/2}\ {\rm erf}(\sqrt{2}/\sigma)$ with $\rm erf$} the error function. For antineutrinos, we assume their angular distribution follows $\bar{g}(v_{z},\sigma)=\alpha g(v_{z},\sigma)$ with $\alpha=n_{\bar{\nu}_{e}}/n_{\nu_{e}}$ being the neutrino-antineutrino asymmetry factor. The choice of the same width parameter $\sigma$ for neutrinos and antineutrinos is to avoid appearance of ELN crossings that can trigger FFCs. In the absence of an ELN angular crossing, no unstable modes exist in the limit $\omega\rightarrow 0$, and any instability is due to SFC. In the case of SFC, a crossing in the energy spectrum is required for modes to become unstable. In a single-energy setup, this is typically achieved by treating antineutrinos as having “negative” energy ($\omega<0$), since they pick up a minus sign in $H_{\rm vac}$ compared to neutrinos ($\omega>0$), thus guaranteeing a crossing in the energy spectra.
@@
isotropic, NO
isotropic, IO
anisotropic, NO
anisotropic, IO
Figure 4: Distributions (upper), absolute values of the eigenvectors $\vec{V}$ (middle), and a simulation snapshot of $\langle P_{3}\rangle$ and $\langle\bar{P}_{3}\rangle$ in linear regime (bottom). The columns correspond to isotropic angular distributions ($\sigma\rightarrow\infty$, $k=0$) for NO and IO, and anisotropic angular distributions ($\sigma=0.6$, $k=0.075$) for NO and IO, respectively. We keep the parameters $\alpha=0.85$, and $|\omega|=0.01$ fixed. The normalization is such that $\sum|{\vec{V}}|^{2}=1$. In the upper panels, the color distinguishes the angular bins from 1-1 (red) to 11 (purple), while thick and thin markers correspond to $\nu_{e}$ and $\bar{\nu}_{e}$, respectively. The bottom panels show the angular distributions of the averaged $P_{3}$ and $\bar{P}_{3}$ in the linear regime ($t=60$ for NO, and $t=20$ for IO). The angular modes at which $P_{3},\bar{P}_{3}$ are closest to 1 correspond to the locations where $|{\vec{V}}|$ has a minimum. The shape of the angular distribution in the linear regime is described by the shape of $|{\vec{V}}|$.
@@ IV Linear Stability Analysis
$\rho$ $=$ $\begin{pmatrix}\rho_{ee}(t_{0})-\Delta(t)&\rho_{ex}(t)\\ \rho_{ex}^{*}(t)&\rho_{xx}(t_{0})+\Delta(t)\end{pmatrix}\$, (13)
$\bar{\rho}$ $=$ $\begin{pmatrix}\bar{\rho}_{ee}(t_{0})-\bar{\Delta}(t)&\bar{\rho}_{ex}(t)\\ \bar{\rho}_{ex}^{*}(t)&\bar{\rho}_{xx}(t_{0})+\bar{\Delta}(t)\end{pmatrix}\$, (14)
@@
[…]
In Fig. 2, we show the dispersion relation of a system with $\omega=\pm 0.01$ and (anti)neutrino angular distributions defined by Eqs. 8 and 10 with $\alpha=0.850$ and $\sigma=0.6$. The NO case (blue) generically shows smaller growth rates than the IO case (red) across the $k$-domain. Moreover, the IO case has a wider range of instability in $k$ than the NO case. Both mass orderings have a global maximum around the $k=0.075$ mode which is marked by the dotted line. Additionally, one can see that the IO case is approximately dominated by the $k=0$ mode, while the NO case has a subdominant homogeneous mode and is thus driven by $k\neq 0$ modes. To paint a more general picture of the instability region, in Fig. 3 we show the growth rates of the parameter space comprised by $0.8\leq\alpha\leq 1.2$ and $0.02\leq\omega\leq 0.02$. In the top panel of Fig. 3, one can see that for the NO case, the systems are more strongly dominated by inhomogeneous ($k\neq 0$) modes when $\alpha$ deviates from 1 as previously discussed. Moreover, the growth rates are approximately symmetric around $\alpha=1$, e.g., the growth rates of mode $k$ for $\alpha=0.9$ is roughly the same as that of mode k-k for $\alpha=1.1$. The bottom panel of Fig. 3 shows the instability region for the dominant $k$ mode for any given pair of parameters $(\omega,\alpha)$. One can see that the $\omega<0$ (IO) quadrant has generally larger growth rates that span a wider range of values of $\alpha$ than $\omega>0$ (NO).
@@ V Flavor Equilibration of SN Neutrinos Induced by Slow Modes
Flavor equilibration for $\boldsymbol{\alpha\leq 1}$.—As illustrated in Fig. 5, the averaged survival probabilities for neutrinos and antineutrinos reach equilibration after a few $\omega^{-1}$ in both neutrino mass orderings. While the antineutrinos approach a state closer to equipartition, the neutrinos settle at a value consistent with the conservation of the total (E–X)LN (which is guaranteed to a good degree given the suppressed effective vacuum mixing angle; see Appendix A). Due to the approximate conservation of (E–X)LN, the final flavor conversion of the neutrinos is related to that of the antineutrinos through
@@
Figure 9: Time evolution of the $P_{3}(t,z_{,}v_{z})$ component of the neutrino polarization vector in the domain box spanned by the spatial variable $50\leq z\leq 50$ (horizontal axis) and the angular variable $1\leq v_{z}\leq 1$ (vertical axis). The evolution corresponds to the benchmark system with $\alpha=0.925$ and $\sigma=0.6$. The panels in the left column correspond to snapshots of the flavor evolution for the $\omega=0.01$ (NO) case, while the right panels to the $\omega=-0.01$ (IO) case.
Dependence on mass ordering and role of instabilities. —In almost all cases, antineutrinos in the IO tend to evolve toward a state closer to perfect equipartition. While there is no obvious explanation for this difference, we speculate that this behavior is associated with the distinct instability characteristics of the neutrino gas in the linear regime for NO and IO. In particular, as illustrated in Fig. 2, in IO, a wider range of Fourier modes is unstable in the linear regime (and with higher growth rate), implying that the initial evolution of inhomogeneity-induced decoherence could proceed more efficiently in IO, with the homogeneous mode ($k=0$) being almost the dominantly unstable mode. This fundamental difference between NO and IO is also associated with the different shapes of the eigenvectors of the unstable modes, as illustrated in Fig. 4. While in the IO case, all angle bins are almost equally affected by instabilities, in the NO case, the impact varies depending on the angle.
@@
Impact of inhomogeneities and comparison to previous studies.—While near flavor equipartition was also observed to occur due to the multi-angle-induced decoherence previously reported in 1D models for $\alpha$ values very close to 1 for IO [93], our key finding is that in the presence of inhomogeneities, this becomes a more generic outcome for the neutrino gas. As shown in Fig. 10, the neutrino gas reaches near flavor equipartition for $\alpha=1$ cases without advection in IO (consistent with what was observed in homogeneous models). Near equipartition for all other cases occurs only when advection is included in the model. This implies that the equilibration unearthed in this study can be considered as an inhomogeneity-induced decoherence rather than a multi-angle-induced decoherence discovered previously in the literature.
@@
Flavor equilibration for $\boldsymbol{\alpha>1}$ and angular distribution features.—So far, we have only discussed cases where $\alpha\leq 1$. A very similar flavor equilibration outcome can occur for $\alpha>1$, as shown in Fig. 11. However, there is an important yet predictable difference in this scenario: the roles of neutrinos and antineutrinos are reversed. Specifically, for $\alpha>1$, it is the neutrinos that reach near flavor equipartition, while the antineutrinos adjust to a value that ensures the approximate (E-X)LN conservation. In addition and similar to $\alpha<1$, neutrinos reach a state closer to equipartition for IO.